# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v5

@@ 1 Introduction
Pricing plays a central role in competitive markets, where firms continuously adjust prices in response to demand fluctuations and rival strategies. A major challenge in competition-aware pricing lies in inferring rivals’ pricing behavior from limited observations (li2024lego). Firms cannot easily estimate price sensitivity through controlled experiments since competitors do not coordinate to hold prices constant while one firm tests different price values. Although existing sequential pricing algorithms yield low regret and converge toward a Nash Equilibrium (NE), they often rely on a linear demand framework (kirman1975learning; li2024lego), or nonlinear approaches restricted to a fixed parametric family (goyal2023learning), limiting applicability. Because nonlinear demand better reflects reality (gallego2006price; wan2022nonlinear; wan2023conditions), we adopt a flexible semiparametric model with unknown parametric and nonparametric components.
Over $T$ periods, $N$ sellers set prices simultaneously; each seller’s demand depends on both their own and rivals’ prices, an effect amplified beyond linearity. Sellers observe competitors’ prices but not competitors’ realized demand. For analysis, we assume that all sellers use the same algorithm – a realistic simplification: for instance, PricingServiceAI2025 describes a pricing service used by numerous local hotels in Colorado. These hotels, while independently operated, utilize the same class of algorithmic tools provided by the platform to set their prices. Importantly, although they observe competitors’ prices, they do not share underlying demand information with one another, as we assume in this work.
@@
In this paper, we assume that the mean demand for a seller $i\in[N]=\{1,2,\dots,N\}$, given the price vector $(p_{i},\mathbf{p}_{-i})$ (where $\mathbf{p}_{-i}$ denots the competitors’ prices), follows a single index model of the form $\mathbb{E}[y_{i}|p_{i},\mathbf{p}_{-i}]=\psi_{i}(-\beta_{i}p_{i}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle)$, where the demand link function $\psi_{i}$ is assumed to be both monotone and $s$-concave ($\langle\cdot,\cdot\rangle$ denotes the euclidean inner product) and $\beta_{i}>0$, which generalizes the linear demand setting in li2024lego, where $\psi_{i}$ is linear (that is monotone and log-concave, i.e. $s$-concave with $s=0$). Below, we summarize the motivation and relevance of these shape constraints, whose importance will become further evident throughout the paper.
Monotonicity aligns with economic intuition, where the standard form of demand is decreasing in the seller’s own price (li2024lego; friesz2012competitive; kirman1975learning), and consequently, since $\beta_{i}>0$, the functions $\psi_{i}$ must be increasing to preserve decreasing demand with respect to $p_{i}$. Indeed, in this case, $\partial_{p_{i}}\mathbb{E}[y_{i}\mid p_{i},\mathbf{p}_{-i}]<0$.
The $s$-concavity assumption of $\psi_{i}$ also arises naturally as a generalization of the commonly used log-concavity (recovered when $s=0$), and plays a crucial role in guaranteeing convergence analysis towards equilibrium (see Section˜3.1). We emphasize that $s$-concavity, being a higher-order smoothness constraint than monotonicity, drives the convergence rate results.
@@ 2 Related Literature and Contributions
Classic price competition with known demand goes back to cournot1838recherches; bertrand1883review, with later variants including multinomial logit (gallego2006price; aksoy2013price; gallego2014multiproduct), fixed-point methods for mixed logit (morrow2011fixed), and multi-epoch competition (gallego2014dynamic; federgruen2015multi; chen2021duopoly). However, over the last two decades, demand learning for competitive pricing has advanced substantially.
In this paper, we study sequential price competition with demand learning. Early work (kirman1975learning) analyzed symmetric duopolies with linear demand. li2024lego achieved optimal $\sqrt{T}$ regret for asymmetric linear demand, $\lambda_{i}(\mathbf{p})=\alpha_{i}-\beta_{i}p_{i}+\sum_{j\neq i}\gamma_{ij}p_{j}$, with unknown sensitivities. However, nonlinear demand is more realistic and widely used (gallego2006price; wan2022nonlinear; wan2023conditions). For instance, gallego2006price study $\lambda_{i}(\mathbf{p})=a_{i}(p_{i})/\big(\sum_{j}a_{j}(p_{j})+\kappa\big)$ for known increasing $a_{i}$ and $\kappa\in(0,1]$. In contrast, we analyze an unknown monotone single-index model
@@
where $\bm{\theta}_{i}$ has $\beta_{i}$ in coordinate $i$ and $\gamma_{ij}$ elsewhere, and both $\bm{\theta}_{i}$ and $\psi_{i}$ are unknown. This strictly generalizes li2024lego: setting $\psi_{i}(u)=u+\alpha_{i}$ recovers their linear model.
@@
By Equation˜1, we have $N$ single-index models $\mathbb{E}(y_{i}\mid\mathbf{p})=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle),\ i\in\mathcal{N}=\{1,2,\dots,N\},$ with unknown $\bm{\theta}_{i}\in\mathbb{R}^{N}$ and unknown links $\psi_{i}$. We assume $\psi_{i}$ is nondecreasing, which makes demand $\lambda_{i}$ nonincreasing in its own price $p_{i}$, a standard assumption in the economics literature (birge2024interfere; li2024lego). Monotonicity also permits fully data-driven, tuning-free nonparametric estimators (balabdaoui2019least). The monotone SIM is identifiable under mild conditions: $\|\bm{\theta}_{i}\|_{2}=1$, and $\mathbf{p}$ has a strictly positive density on its domain (see balabdaoui2019least). balabdaoui2019least also propose a joint estimator $(\widehat{\bm{\theta}}_{i,n},\widehat{\psi}_{i,n})$ with $L^{2}$ error of order $n^{-1/3}$ for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$, and study the normalized linear estimator
@@
which remains consistent under elliptically symmetric $\mathbf{p}^{(t)}$ (brillinger2012generalized). In this work, we also assume that $\mathbf{p}^{(t)}$ follows an elliptically symmetric distribution during the exploration phase. However, for our purpose, an $L^{2}$ convergence alone for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$ is insufficient to establish a regret bound, but a uniform (supremum-norm) rate suffices and is derived in this paper.
@@
We propose a sequential price competition online algorithm that provides sublinear regret. The first step to achieve this goal is to establish the existence and uniqueness of a Nash Equilibrium $\mathbf{p^{*}}$, which is the fixed point of an operator $\bm{\Gamma}$ (see Section˜3.1 for more details). A necessary condition for its existence is that each seller’s virtual valuation
@@
is increasing with derivative bounded below: $\varphi_{i}^{\prime}(u)\geq c_{i}$ for some $c_{i}>0$ for all $i\in\mathcal{N}$. This mirrors the $N=1$ literature (monopolistic setting) such as fan2024policy; chen2018robust; cole2014sample; golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic, which often assume log-concavity of $\psi_{i}$ and $1-\psi_{i}$ (implying $c_{i}\geq 1$).
Our key observation links the NE and shape constraints as follows: $\varphi_{i}^{\prime}(u)\geq c_{i}$ iff $\psi_{i}$ is $(c_{i}-1)$-concave (Proposition˜3.5; see the section below for definition of $s$-concavity), with log-concavity recovered as a special case when $c_{i}=1$. This connection allows us to estimate $\psi_{i}$ via nonparametric least squares under $s$-concavity – yielding a fully data-driven, tuning-free estimation procedure.
@@
A technical contribution of this work is the study of the uniform rate of convergence of the least-squares estimator of a unidimensional $s$-concave regression function. As defined in han2016approximation, a unidimensional function $\psi:\mathcal{U}\rightarrow(0,\infty)$, $\mathcal{U}\subset\mathbb{R}$, is said to be $s$-concave for some $s\in\mathbb{R}$, and we write $\psi\in\mathcal{F}_{s}(\mathcal{U})$ if $\psi\left((1-\lambda)u_{0}+\lambda u_{1}\right)\geq M_{s}\left(\psi\left(u_{0}\right),\psi\left(u_{1}\right);\lambda\right)$, for all $u_{0},u_{1}\in\mathcal{U}$ and $\lambda\in(0,1)$, where
@@
This notion generalizes concavity ($s=1$) and log-concavity (which holds for $s=0$, in the sense that $\lim_{s\rightarrow 0}M_{s}(y_{0},y_{1};\lambda)=M_{0}(y_{0},y_{1};\lambda)$). These classes are nested since $\mathcal{F}_{s}(\mathcal{U})\subset\mathcal{F}_{0}(\mathcal{U})\subset\mathcal{F}_{r}(\mathcal{U})$, if $\infty<r<0<s<\infty$. The class of log-concave densities has been extensively studied: see bobkov2011concentration; dumbgen2009maximum; cule2010theoretical; borzadaran2011log; bagnoli2006log; while han2016approximation; doss2016global; chandrasekaran2009sampling; koenker2010quasi deal with general $s$-concavity. It is easy to see that such functions $\psi$ have the form $\psi=(\phi_{+})^{1/s}$ for some concave function $\phi$ if $s>0$, where $x_{+}=\max\{0,x\}$, $\psi=e^{\phi}$ for some concave function $\phi$ if $s=0$, and $\psi=(\phi_{+})^{1/s}$ for some convex function $\phi$ if $s<0$. Then, $\psi$ has the following representation: $\psi=h_{s}\circ\phi$ where $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is concave and
@@
Different from previous results, we study the convergence rate for the more general class $\mathcal{F}_{h}$ (see Appendix˜H for details) that contains functions of the form $\psi=h\circ\phi:\mathcal{U}\to\mathbb{R}$, where $h:\mathbb{R}\rightarrow(0,\infty)$ is a known increasing function in $\mathrm{C}^{2}$ and $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is an unknown concave function. Specifically, we consider i.i.d. observations $(U_{i},Y_{i})\sim(U,Y)$ for $i=1,2,\dots,n$ where $Y_{1},\dots,Y_{n}$ are noisy representations of the mean function $\psi_{0}(u)=\mathbb{E}\left(Y|U=u\right)=h(\phi_{0}(u))$ and $U_{1}\leq U_{2}\leq\dots\leq U_{n}$ are contained in $\mathcal{U}$. We study the uniform convergence rate of $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$, where $\widehat{\phi}_{n}$ is the LSE
@@ 2.1 Summary of Key Contributions
A novel semiparametric pricing policy for nonlinear mean demand. We extend the standard approach to estimating the mean demand function of a firm $\lambda_{i}(\cdot)$ by introducing a monotone single index model $\lambda_{i}(\mathbf{p})\;=\;\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle)$, where $\psi_{i}$ is increasing and $s_{i}$-concave for some $s_{i}>-1$, providing substantially more flexibility than previous parametric models (li2024lego; kachani2007modeling; gallego2006price).
$s$-concave mean demand functions. In many existing works, necessary conditions for the existence of the NE are derived by assuming that the virtual valuation function defined in Equation˜3 is increasing, often by invoking log-concavity of the mean demand function. We show how all such assumptions can be cast under the more general framework of $s$-concavity. This reformulation allows the development of a fully data-driven, tuning-parameter-free algorithm using shape constraints.
Regret upper bound and convergence to equilibrium. We establish an upper bound on the total expected regret and analyze the convergence to the NE for a general exploration length of order $\tau\propto T^{\xi}$ for $\xi\in(0,1)$ (Theorem˜5.4). Our results reveal the existence of an optimal choice of $\xi$ that minimizes the total expected regret for each seller, leading to a regret of order $\widetilde{\mathcal{O}}(N^{3/2}T^{5/7})$, where $\widetilde{\mathcal{O}}$ excludes log factors (Remark˜5.6). Moreover, we show that by the end of the selling horizon, the joint prices set by sellers converge to Nash equilibrium prices at a rate of $\widetilde{\mathcal{O}}(N^{3/4}T^{-1/7})$.
Concentration inequality for $s$-concave regression functions in the supremum norm. Our work involves establishing a concentration inequality for the nonparametric LSE, under the supremum norm, for a large class of shape constraints that includes $s$-concavity (Appendix˜H). As a minor contribution, we derive a concentration inequality for the parametric component $\bm{\theta}_{i}$ of the monotone single index model (Proposition˜D.3), while previous results show convergence in probability or distribution (balabdaoui2019least).

### v6

@@ 1 Introduction
Pricing plays a central role in competitive markets, where firms continuously adjust prices in response to demand fluctuations and rival strategies. A major challenge in competition-aware pricing lies in inferring rivals’ pricing behavior from limited observations (Li et al., 2024). Firms cannot easily estimate price sensitivity through controlled experiments since competitors do not coordinate to hold prices constant while one firm tests different price values. Although existing sequential pricing algorithms yield low regret and converge toward a Nash Equilibrium (NE), they often rely on a linear demand framework (Kirman, 1975; Li et al., 2024), or nonlinear approaches restricted to a fixed parametric family (Goyal et al., 2023), limiting applicability. Because nonlinear demand better reflects reality (Gallego et al., 2006; Wan et al., 2022; 2023), we adopt a flexible semiparametric model with unknown parametric and nonparametric components.
Over $T$ periods, $N$ sellers set prices simultaneously; each seller’s demand depends on both their own and rivals’ prices, an effect amplified beyond linearity. Sellers observe competitors’ prices but not competitors’ realized demand. For analysis, we assume that all sellers use the same algorithm – a realistic simplification: for instance, 43 describes a pricing service used by numerous local hotels in Colorado. These hotels, while independently operated, utilize the same class of algorithmic tools provided by the platform to set their prices. Importantly, although they observe competitors’ prices, they do not share underlying demand information with one another, as we assume in this work.
@@
In this paper, we assume that the mean demand for a seller $i\in[N]=\{1,2,\dots,N\}$, given the price vector $(p_{i},\mathbf{p}_{-i})$ (where $\mathbf{p}_{-i}$ denots the competitors’ prices), follows a single index model of the form $\mathbb{E}[y_{i}|p_{i},\mathbf{p}_{-i}]=\psi_{i}(-\beta_{i}p_{i}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle)$, where the demand link function $\psi_{i}$ is assumed to be both monotone and $s$-concave ($\langle\cdot,\cdot\rangle$ denotes the euclidean inner product) and $\beta_{i}>0$, which generalizes the linear demand setting in Li et al. (2024), where $\psi_{i}$ is linear (that is monotone and log-concave, i.e. $s$-concave with $s=0$). Below, we summarize the motivation and relevance of these shape constraints, whose importance will become further evident throughout the paper.
Monotonicity aligns with economic intuition, where the standard form of demand is decreasing in the seller’s own price (Li et al., 2024; Friesz et al., 2012; Kirman, 1975), and consequently, since $\beta_{i}>0$, the functions $\psi_{i}$ must be increasing to preserve decreasing demand with respect to $p_{i}$. Indeed, in this case, $\partial_{p_{i}}\mathbb{E}[y_{i}\mid p_{i},\mathbf{p}_{-i}]<0$.
The $s$-concavity assumption of $\psi_{i}$ also arises naturally as a generalization of the commonly used log-concavity (recovered when $s=0$), and plays a crucial role in guaranteeing convergence analysis towards equilibrium (see Section 3.1). We emphasize that $s$-concavity, being a higher-order smoothness constraint than monotonicity, drives the convergence rate results.
@@ 2 Related Literature and Contributions
Classic price competition with known demand goes back to Cournot (1838); Bertrand (1883), with later variants including multinomial logit (Gallego et al., 2006; Aksoy-Pierson et al., 2013; Gallego and Wang, 2014), fixed-point methods for mixed logit (Morrow and Skerlos, 2011), and multi-epoch competition (Gallego and Hu, 2014; Federgruen and Hu, 2015; Chen and Chen, 2021). However, over the last two decades, demand learning for competitive pricing has advanced substantially.
In this paper, we study sequential price competition with demand learning. Early work (Kirman, 1975) analyzed symmetric duopolies with linear demand. Li et al. (2024) achieved optimal $\sqrt{T}$ regret for asymmetric linear demand, $\lambda_{i}(\mathbf{p})=\alpha_{i}-\beta_{i}p_{i}+\sum_{j\neq i}\gamma_{ij}p_{j}$, with unknown sensitivities. However, nonlinear demand is more realistic and widely used (Gallego et al., 2006; Wan et al., 2022; 2023). For instance, Gallego et al. (2006) study $\lambda_{i}(\mathbf{p})=a_{i}(p_{i})/\big(\sum_{j}a_{j}(p_{j})+\kappa\big)$ for known increasing $a_{i}$ and $\kappa\in(0,1]$. In contrast, we analyze an unknown monotone single-index model
@@
where $\bm{\theta}_{i}$ has $\beta_{i}$ in coordinate $i$ and $\gamma_{ij}$ elsewhere, and both $\bm{\theta}_{i}$ and $\psi_{i}$ are unknown. This strictly generalizes Li et al. (2024): setting $\psi_{i}(u)=u+\alpha_{i}$ recovers their linear model.
@@
By Equation 1, we have $N$ single-index models $\mathbb{E}(y_{i}\mid\mathbf{p})=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle),\ i\in\mathcal{N}=\{1,2,\dots,N\},$ with unknown $\bm{\theta}_{i}\in\mathbb{R}^{N}$ and unknown links $\psi_{i}$. We assume $\psi_{i}$ is nondecreasing, which makes demand $\lambda_{i}$ nonincreasing in its own price $p_{i}$, a standard assumption in the economics literature (Birge et al., 2024; Li et al., 2024). Monotonicity also permits fully data-driven, tuning-free nonparametric estimators (Balabdaoui et al., 2019). The monotone SIM is identifiable under mild conditions: $\|\bm{\theta}_{i}\|_{2}=1$, and $\mathbf{p}$ has a strictly positive density on its domain (see Balabdaoui et al. (2019, Prop. 5.1)). Balabdaoui et al. (2019) also propose a joint estimator $(\widehat{\bm{\theta}}_{i,n},\widehat{\psi}_{i,n})$ with $L^{2}$ error of order $n^{-1/3}$ for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$, and study the normalized linear estimator
@@
which remains consistent under elliptically symmetric $\mathbf{p}^{(t)}$ (Brillinger, 2012). In this work, we also assume that $\mathbf{p}^{(t)}$ follows an elliptically symmetric distribution during the exploration phase. However, for our purpose, an $L^{2}$ convergence alone for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$ is insufficient to establish a regret bound, but a uniform (supremum-norm) rate suffices and is derived in this paper.
@@
We propose a sequential price competition online algorithm that provides sublinear regret. The first step to achieve this goal is to establish the existence and uniqueness of a Nash Equilibrium $\mathbf{p^{*}}$, which is the fixed point of an operator $\bm{\Gamma}$ (see Section 3.1 for more details). A necessary condition for its existence is that each seller’s virtual valuation
@@
is increasing with derivative bounded below: $\varphi_{i}^{\prime}(u)\geq c_{i}$ for some $c_{i}>0$ for all $i\in\mathcal{N}$. This mirrors the $N=1$ literature (monopolistic setting) such as Fan et al. (2024); Chen and Farias (2018); Cole and Roughgarden (2014); Golrezaei et al. (2019); Javanmard (2017); Javanmard and Nazerzadeh (2019), which often assume log-concavity of $\psi_{i}$ and $1-\psi_{i}$ (implying $c_{i}\geq 1$).
Our key observation links the NE and shape constraints as follows: $\varphi_{i}^{\prime}(u)\geq c_{i}$ iff $\psi_{i}$ is $(c_{i}-1)$-concave (Proposition 3.5; see the section below for definition of $s$-concavity), with log-concavity recovered as a special case when $c_{i}=1$. This connection allows us to estimate $\psi_{i}$ via nonparametric least squares under $s$-concavity – yielding a fully data-driven, tuning-free estimation procedure.
@@
A technical contribution of this work is the study of the uniform rate of convergence of the least-squares estimator of a unidimensional $s$-concave regression function. As defined in Han and Wellner (2016), a unidimensional function $\psi:\mathcal{U}\rightarrow(0,\infty)$, $\mathcal{U}\subset\mathbb{R}$, is said to be $s$-concave for some $s\in\mathbb{R}$, and we write $\psi\in\mathcal{F}_{s}(\mathcal{U})$ if $\psi\left((1-\lambda)u_{0}+\lambda u_{1}\right)\geq M_{s}\left(\psi\left(u_{0}\right),\psi\left(u_{1}\right);\lambda\right)$, for all $u_{0},u_{1}\in\mathcal{U}$ and $\lambda\in(0,1)$, where
@@
This notion generalizes concavity ($s=1$) and log-concavity (which holds for $s=0$, in the sense that $\lim_{s\rightarrow 0}M_{s}(y_{0},y_{1};\lambda)=M_{0}(y_{0},y_{1};\lambda)$). These classes are nested since $\mathcal{F}_{s}(\mathcal{U})\subset\mathcal{F}_{0}(\mathcal{U})\subset\mathcal{F}_{r}(\mathcal{U})$, if $\infty<r<0<s<\infty$. The class of log-concave densities has been extensively studied: see Bobkov and Madiman (2011); Dümbgen and Rufibach (2009); Cule and Samworth (2010); Borzadaran and Borzadaran (2011); Bagnoli and Bergstrom (2006); while Han and Wellner (2016); Doss and Wellner (2016); Chandrasekaran et al. (2009); Koenker and Mizera (2010) deal with general $s$-concavity. It is easy to see that such functions $\psi$ have the form $\psi=(\phi_{+})^{1/s}$ for some concave function $\phi$ if $s>0$, where $x_{+}=\max\{0,x\}$, $\psi=e^{\phi}$ for some concave function $\phi$ if $s=0$, and $\psi=(\phi_{+})^{1/s}$ for some convex function $\phi$ if $s<0$. Then, $\psi$ has the following representation: $\psi=h_{s}\circ\phi$ where $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is concave and
@@
Different from previous results, we study the convergence rate for the more general class $\mathcal{F}_{h}$ (see Appendix H for details) that contains functions of the form $\psi=h\circ\phi:\mathcal{U}\to\mathbb{R}$, where $h:\mathbb{R}\rightarrow(0,\infty)$ is a known increasing function in $\mathrm{C}^{2}$ and $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is an unknown concave function. Specifically, we consider i.i.d. observations $(U_{i},Y_{i})\sim(U,Y)$ for $i=1,2,\dots,n$ where $Y_{1},\dots,Y_{n}$ are noisy representations of the mean function $\psi_{0}(u)=\mathbb{E}\left(Y|U=u\right)=h(\phi_{0}(u))$ and $U_{1}\leq U_{2}\leq\dots\leq U_{n}$ are contained in $\mathcal{U}$. We study the uniform convergence rate of $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$, where $\widehat{\phi}_{n}$ is the LSE
@@ 2.1 Summary of Key Contributions
A novel semiparametric pricing policy for nonlinear mean demand. We extend the standard approach to estimating the mean demand function of a firm $\lambda_{i}(\cdot)$ by introducing a monotone single index model $\lambda_{i}(\mathbf{p})\;=\;\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle)$, where $\psi_{i}$ is increasing and $s_{i}$-concave for some $s_{i}>-1$, providing substantially more flexibility than previous parametric models (Li et al., 2024; Kachani et al., 2007; Gallego et al., 2006).
$s$-concave mean demand functions. In many existing works, necessary conditions for the existence of the NE are derived by assuming that the virtual valuation function defined in Equation 3 is increasing, often by invoking log-concavity of the mean demand function. We show how all such assumptions can be cast under the more general framework of $s$-concavity. This reformulation allows the development of a fully data-driven, tuning-parameter-free algorithm using shape constraints.
Regret upper bound and convergence to equilibrium. We establish an upper bound on the total expected regret and analyze the convergence to the NE for a general exploration length of order $\tau\propto T^{\xi}$ for $\xi\in(0,1)$ (Theorem 5.4). Our results reveal the existence of an optimal choice of $\xi$ that minimizes the total expected regret for each seller, leading to a regret of order $\widetilde{\mathcal{O}}(N^{3/2}T^{5/7})$, where $\widetilde{\mathcal{O}}$ excludes log factors (Remark 5.6). Moreover, we show that by the end of the selling horizon, the joint prices set by sellers converge to Nash equilibrium prices at a rate of $\widetilde{\mathcal{O}}(N^{3/4}T^{-1/7})$.
Concentration inequality for $s$-concave regression functions in the supremum norm. Our work involves establishing a concentration inequality for the nonparametric LSE, under the supremum norm, for a large class of shape constraints that includes $s$-concavity (Appendix H). As a minor contribution, we derive a concentration inequality for the parametric component $\bm{\theta}_{i}$ of the monotone single index model (Proposition D.3), while previous results show convergence in probability or distribution (Balabdaoui et al., 2019).